A simple random sample of 2020 pages from a dictionary is obtained. The numbers of words defined on those pages are found, with the results nequals=2020 , x overbarxequals=55.255.2 words, sequals=16.516.5 words. Given that this dictionary has 14341434 pages with defined words, the claim that there are more than 70,000 defined words is equivalent to the claim that the mean number of words per page is greater than 48.848.8 words. Use a 0.010.01 significance level to test the claim that the mean number of words per page is greater than 48.848.8 words. What does the result suggest about the claim that there are more than 70,000 defined words? Identify the null and alternative hypotheses, test statistic, P-value, and state the final conclusion that addresses the original claim. Assume that the population is normally distributed
Answer #1

H0 ; mu = 48.8
Ha: mu > 48.8

test statistics:

xbar = 55.2 , s = 16.5 , n = 20

t = (xbar -mu)/(s/sqrt(n))
= ( 55.2 - 48.8)/(16.5/sqrt(20))
= 1.73

p value = 1- P(t < 1.73) with df = 19

P value = 0.0499


As p value > 0.01 fail to reject H0

There is not sufficient evidence to support the claim that the mean number of words per page is greater than 48.8